OpenStudy (anonymous):

Find the annual growth rate if a population has grown exponentially from 40 million to 50 million in the last 20 years. At this rate, what will be the population 10 years from now?

OpenStudy (anonymous):

Okay, what can you do on your own?

OpenStudy (anonymous):

We have the formula: \[ a(t) = a_0r^t \]

OpenStudy (anonymous):

We can let \(a_0=50\) in millions, and say that \(a(-10) = 40\). Then we just want \(a(20)\).

OpenStudy (anonymous):

\[ 40 = 50r^{-10} \]Can you solve for \(r\)?

OpenStudy (anonymous):

how come it's a negative 10? 5/4r^-10?

OpenStudy (anonymous):

We are saying it was initially at 50, and 10 years in the past it was 40.

OpenStudy (anonymous):

oh okay i understand

OpenStudy (anonymous):

to solve for r do i just divide 50 to both sides?

OpenStudy (anonymous):

That's a start.

OpenStudy (anonymous):

5/4?

OpenStudy (anonymous):

No, not quite.

OpenStudy (anonymous):

can you help go through it then?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

\[ \frac{4}{5} =r^{-10} \\ \frac54=r^{10}\\ r = \sqrt[10]{\frac 54} \]

OpenStudy (anonymous):

Actually, I have a better idea. Let's just say that \(t\) is for every \(10\) years, instead of for every \(1\) year.

OpenStudy (anonymous):

If we do that, then \[ r = \frac 54 \]

OpenStudy (anonymous):

And we just want \(a(2)\) where: \[ a(t) = 50\left(\frac 54\right)^t \]

OpenStudy (anonymous):

so… 2=50(5/4)^t ?

OpenStudy (anonymous):

No, not \(a(t)=2\). I said \(a(2)\). \(t=2\).

OpenStudy (anonymous):

oh so just 50(5/4)^2 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Wait, I am wrong..

OpenStudy (anonymous):

whats wrong

OpenStudy (anonymous):

It should be \(a(1/2)\).

OpenStudy (anonymous):

I said that \(t\) will be every ten years. But it really was every 20 years. So ten years from now is \(1/2\)

OpenStudy (anonymous):

55.90 for the answer?

OpenStudy (anonymous):

and how did you get 1/2?

OpenStudy (anonymous):

@wio